OpenStudy (anonymous):

what is 216 to the power of 4/3?

OpenStudy (anonymous):

1296, it's the same as \[\sqrt[3]{216}\] to the power of 4

OpenStudy (uri):

\[216 ^{4/3} ?\]

OpenStudy (anonymous):

I.e 6^4

OpenStudy (uri):

That's the question right?

OpenStudy (uri):

You have a calculator? are you allowed to use it @corybory

OpenStudy (anonymous):

\[x^{a/b} = (\sqrt[b]{x})^a\]

OpenStudy (anonymous):

As 216 is a perfect cube, this method works here, otherwise you would need a calculator

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

No problem, need help with anything else?

OpenStudy (anonymous):

Express a^14/3 b^ 16/3 C^ 5/3 in simplified radical form.